# Human-AI Collaborative Multi-modal Multi-rater Learning for Endometriosis Diagnosis

### v2

@@ 1 Introduction
The T1 and T2 modalities of Magnetic Resonance Imaging (MRI) are among the most recommended medical imaging methods for diagnosing endometriosis given their effectiveness to visualize many signs of the condition. One of the most important signs associated with the condition is the Pouch of Douglas (POD) Obliteration [7, 8]. Developing an AI model capable of classifying POD obliteration has the potential to facilitate the widespread adoption of imaging-based diagnosis and enhance diagnosis accuracy and consistency. However, training such a model relies on acquiring precise POD Obliteration annotations from T1/T2 MRIs, which is a challenging task because even experienced clinicians may lack certainty regarding the presence of the sign. In fact, the uncertainty in the manual POD obliteration classification from T1/T2 MRI is remarkably low, with only 61.4% to 71.9% accuracy [9, 10]. Nevertheless, there have been some attempts at training such multi-modal MRI AI models for classifying POD obliteration, such as Zhang et al.’s [11] method to distill the POD obliteration classification knowledge from ultrasound to MRI, or Butler et al. [8]’s self-supervised pre-training for multi-modal POD obliteration classification. However, none of these methods have been tested against ground truth labels obtained from surgery reports, so it is hard to assess if their published accuracy results are competitive with purely manual classifications. Therefore, a major research question in this problem is if it is possible to design innovative training and testing methodologies that can lead to highly accurate POD obliteration classification results.
@@ 2.4 Imaging-based Endometriosis Detection
One crucial indicator to detect endometriosis is the obliteration of the Pouch of Douglas (POD) [7, 8]. However, the development of an AI model that can classify such indicator hinges on the availability of precise POD obliteration annotations from T1/T2 MRIs, a task that is challenging because even experienced clinicians often face uncertainty in identifying this sign [56, 57, 58, 59]. Despite these challenges, there have been some efforts to train multi-modal MRI AI models for POD obliteration classification. For example, Zhang et al. [11] proposed a method to transfer knowledge from ultrasound to MRI for classifying POD obliteration, and Butler et al. [8] explored self-supervised pre-training for multi-modal POD obliteration classification. However, these methods have not been validated against ground-truth labels obtained from surgical reports, making it difficult to determine if their reported accuracy is truly competitive with manual classification.
@@ 3 Methodology
Figure 1: The framework of HAICOMM. The MRI encoders of HAICOMM are: (a) firstly pre-trained with a Masked Autoencoder (MAE) model; then (b) the pseudo clean labels are estimated from the multi-rater learning process; next, (c) the T1 and T2 data, along with the human-produced multi-rater labels are entered into respective feature extraction encoders – the features from three sources are fused for the final prediction. In the figure, “FC” means fully-connected. “Fts” represents features. “Concat” presents concatenation and $\oplus$” denotes the concatenation operation.
@@ 3.1 Multi-modal Encoder Pre-training
where $\|\cdot\|_{2}$ denotes the L2-norm. For the training and evaluation of the human-AI collaborative classifier, we use the feature extractor $g_{\phi^{*}}(.)$, as explained below in Sec. 3.3.
@@ 3.2 Multi-rater Learning
With the multi-rater labels, we first perform majority vote to fetch the most frequently appearing labels per training sample. Let us present the majority vote operation as $h:\mathcal{Y}\to\{0,1\}$. Then, we have the mapping from multi-rater labels to the majority label for each multi-modal sample. As a result, we can form a new dataset with
@@
where $\hat{y}^{(i)}\in\{0,1\}$. With such generated consensus labels from majority vote, we train a classifier $f_{\theta}:\mathcal{X}\times\mathcal{X}\to[0,1]$ that takes the T1 and T2 MRI volumes to optimize a standard binary cross-entropy objective. This classifier, together with $\mathcal{D}_{T}$ will then be used to train the multi-rater learning method CrowdLab [12], denoted by $z:[0,1]\times\mathcal{Y}\to\{0,1\}$, which generates a pseudo clean label for each sample. Formally, this pseudo labeling process can be defined by:
@@ 3.3 Multi-modal Human-AI Collaborative Classification
Given the pseudo clean labels $y^{(i)}$ from (3), the dataset for model training is defined as:
@@
The pre-trained encoder from (1), parameterized by $\phi^{*}$, is utilized to initialize the T1 and T2 MRI feature extractors, respectively defined by $g_{\hat{\phi}}:\mathcal{X}\to\mathcal{F}$ and $g_{\phi^{\prime}}:\mathcal{X}\to\mathcal{F}$, for the classification task. We also need to define a new feature extractor for processing the manual labels with a learnable module defined as $g_{\gamma}:\mathcal{Y}\to\mathcal{F}$. Such manual labels are used in the human-AI collaborative module.
Hence, these feature extractors produce:
$\small\mathbf{F}_{t1}^{(i)}=g_{\hat{\phi}}(\mathbf{x}_{t1}^{(i)}),\ \mathbf{F}_{t2}^{(i)}=g_{\phi^{\prime}}(\mathbf{x}_{t2}^{(i)}),\ \mathbf{F}_{r}^{(i)}=g_{\gamma}(\mathbf{y}^{(i)}).$ (5)
@@
where $\mathbf{p}^{(i)}\in[0,1]^{2}$ denotes the probabilistic prediction, $\oplus$ represents the concatenation operator, and $\sigma$ is the softmax function. Finally, the training of the whole model is performed by minimizing the binary cross-entropy loss, as follows:
@@
where the $p^{(i)}$ is the predicted probability of the positive class of $i$-th sample (i.e., the $2^{nd}$ dimension $\mathbf{p}^{(i)}$ in (6)), and $y^{(i)}$ is the pseudo clean label in $\mathcal{D}^{\prime}_{T}$.
@@ 4.1 Endometriosis Dataset
We first introduce our multi-modal multi-rater dataset annotated with imaging and surgery-based POD obliteration labels for the diagnosis of endometriosis. For the pre-training stage, we collected 5,867 unlabeled T1 MRI volumes and 8,984 unlabeled T2 MRI volumes from patients aged between 18 and 45 years old, where the volumes show female pelvis scans obtained from various MRI machines with varying resolutions. To standardize the data, the volumes were resampled to $1\text{mm}\times 1\text{mm}\times 3\text{mm}$ voxels. Also, we apply 3D contrast-limited adaptive histogram equalization (CLAHE) to enhance image local contrast and refine edge definitions.
For the training of the human-AI collaborative POD obliteration classifier, we collected 82 pairs of T1/T2 MRI volumes with patients aged 18 to 45 years old. These scans were obtained across multiple clinical sites, with each case annotated by three experienced clinicians who work in clinics specialized in the imaging-based diagnosis of endometriosis. The scans show a specific region surrounding the uterus, which is the area where the signs of POD obliteration are more visible. We further collect 30 cases that contain ground truth annotation of POD obliteration from surgical reports. These cases serve as gold standards for testing. We also use CLAHE pre-processing in this dataset.
@@ 4.2 Implementation Details
For model pre-training, the input volumes are either cropped or zero-padded to achieve dimensions of $64\times 128\times 128$ voxels. To maintain consistency with the pre-training dataset, the endometriosis training samples are cropped in the uterine region at the same dimensions as pre-training data, i.e. $64\times 128\times 128$ voxels. In both pre-training and the training of the human-AI collaborative POD obliteration classifier, the multi-modal encoder for each modality is a transformer with 12 blocks. The majority vote classifier has a 3D-ResNet50 as its backbone network [62]. For the human-AI collaborative POD obliteration classifier training, we use 5 epochs for model optimization warming up. AdamW optimizer and base learning rate of 1e-3 with cosine annealing [63] learning rate tuning strategy are adopted. Three multi-rater labels from three different annotators are incorporated into the training process. In the testing phase, the scans are also cropped in the uterine regions and the clinical surgical results serve as the ground truth for accurate evaluation. All of our models are trained for 60 epochs without model selection. Note that the majority voting is only produced for the consensus pseudo label required for the training of the model $f_{\theta}(.)$ to be used by CrowdLab, as explained in Eq. 3. Once the pseudo clean labels are generated by CrowdLab, the majority voting will no longer be needed.
@@ 4.3 Quantitative Evaluation Settings
4.4 Overall Performance
Table 1: The performance comparison of HAICOMM and other models in terms of test accuracy and AUROC (and respective improvements of HAICOMM) with the testing ground-truth labels from surgical reports. “Majority Vote” denotes a purely manual classification using the majority vote of the three annotators. SSR and ProMix w/ “GT1”, “GT2” and “GT3” mean noisy-label learning models trained with labels from annotators #1, #2 and #3, respectively. “CL” denotes noisy-label learning models trained with labels $y^{(i)}$ from Eq. (3) produced by the multi-rater learning method CrowdLab. “HAIC” represents a model that collaborates with the annotators and that is trained with labels from CrowdLab. The best results for each column are in bold.
@@ 4.4 Overall Performance
Human Majority Vote 0.7000±0.0000 14.29% - -
SSR w/ GT1 0.5833±0.0236 37.93% 0.5763±0.0121 53.83%
Noisy SSR w/ GT2 0.6000±0.0408 33.33% 0.5845±0.0112 51.67%
Label SSR w/ GT3 0.5667±0.0157 41.17% 0.5644±0.0572 57.07%
Learning ProMix w/ GT1 0.5333±0.0236 50.01% 0.5381±0.0412 64.75%
ProMix w/ GT2 0.6167±0.0314 29.72% 0.5808±0.0151 52.63%
ProMix w/ GT3 0.5167±0.0471 54.83% 0.5663±0.0245 56.54%
Multi- SSR w/ CL 0.6167±0.0236 29.72% 0.5878±0.0078 50.82%
rater ProMix w/ CL 0.6500±0.0408 23.08% 0.5428±0.0328 63.32%
HAIC SSR w/ HAIC 0.6833±0.0849 17.08% 0.7424±0.0047 19.37%
ProMix w/ HAIC 0.6667±0.0624 20.00% 0.7389±0.0367 19.98%
Ours HAICOMM 0.8000±0.0408 - 0.8865±0.0551 -
The performance results in Table 1 show that the proposed HAICOMM outperforms other competing models by a large margin across the accuracy and AUC measures. Relative improvements vary from 9.10% to 54.83% on accuracy and 19.37% and 64.75% on AUROC. The standard deviation is calculated by inference time bootstrapping.

### v3

@@ 1 Introduction
The T1 and T2 modalities of Magnetic Resonance Imaging (MRI) are among the most recommended medical imaging methods for diagnosing endometriosis given their effectiveness to visualize many signs of the condition. One of the most important signs associated with the condition is the Pouch of Douglas (POD) Obliteration [7, 8]. Developing an AI model capable of classifying POD obliteration has the potential to facilitate the widespread adoption of imaging-based diagnosis and enhance diagnosis accuracy and consistency. However, training such a model relies on acquiring precise POD Obliteration annotations from T1/T2 MRIs, which is a challenging task because even experienced clinicians may lack certainty regarding the presence of the sign. In fact, the uncertainty in the manual POD obliteration classification from T1/T2 MRI is remarkably low, with only 61.4% to 71.9% accuracy [9, 10]. To demonstrate the difficulty in the detection of POD obliteration, we present Figure 1, showing T1-weighted (b) and (d) paired with T2-weighted (a) and (c) MR images in the sagittal plane. The (a) and (b) pair shows a normal POD case, while the (c) and (d) pair shows POD obliteration, with the red arrow pointing to significant adhesion and distortion, indicating the loss of the soft tissue plane separating the uterine fundus from the bowel in the POD. Nevertheless, there have been some attempts to train machine learning models for POD obliteration detection [11, 8]. Despite their reported accuracy, these methods have not been validated against ground truth labels from surgical reports, making it challenging to assess their clinical value in the detection of POD obliteration. Therefore, a major research question in this problem is if it is possible to design innovative training and testing methodologies that can lead to highly accurate and clinically useful POD obliteration classification results.
Figure 1: Paired sagittal MR images, with T1-weighted (b, d) and T2-weighted (a, c) scans. The (a, b) pair shows a normal POD, while the (c, d) pair reveals POD obliteration, highlighted by a red arrow indicating significant adhesion and tissue distortion, demonstrating the loss of the soft tissue plane separating the uterine fundus from the bowel.
@@ 2.4 Imaging-based Endometriosis Detection
One crucial indicator to detect endometriosis is the obliteration of the Pouch of Douglas (POD) [7, 8]. However, the development of an AI model that can classify such indicator hinges on the availability of precise POD obliteration annotations from T1/T2 MRIs, a task that is challenging because even experienced clinicians often face uncertainty in identifying this sign. Despite these challenges, there have been some efforts to train multi-modal MRI AI models for POD obliteration classification. For example, Zhang et al. [11] proposed a method to transfer knowledge from ultrasound to MRI for classifying POD obliteration, and Butler et al. [8] explored self-supervised pre-training for multi-modal POD obliteration classification. However, these methods have not been validated against ground-truth labels obtained from surgical reports, making it difficult to assess their clinical validity.
@@ 3 Methodology
Figure 2: The framework of HAICOMM. The MRI encoders of HAICOMM are: (a) firstly pre-trained with a Masked Autoencoder (MAE) model; then (b) the pseudo clean labels are estimated from the multi-rater learning process; next, (c) the T1 and T2 data, along with the human-produced multi-rater labels are entered into respective feature extraction encoders – the features from three sources are fused for the final prediction. In the figure, “FC” means fully-connected, “Fts” represents features, “Concat” denotes concatenation, and $\oplus$” is the concatenation operation.
@@ 3.1 Multi-modal Encoder Pre-training
where $\|\cdot\|_{2}$ denotes the L2-norm. The optimization in (1) jointly learns the encoder parameters $\phi$ and the decoder parameters $\psi$ to minimize the reconstruction error of the masked patches across all training samples. For the subsequent training and evaluation of the human-AI collaborative classifier, we adopt the pre-trained feature extractor $g_{\phi^{*}}(.)$, as explained below in Sec. 3.3.
@@ 3.2 Multi-rater Learning
With the multi-rater labels, we first perform majority vote to fetch the most frequently appearing labels per training sample. Let us present the majority vote operation as $h:\mathcal{Y}\to\{0,1\}$. Then, we have the mapping from multi-rater labels to the majority label for each multi-modal sample, forming the following majority voting dataset:
@@
where $\hat{y}^{(i)}\in\{0,1\}$. With such generated consensus labels from majority vote, we train a classifier $f_{\theta}:\mathcal{X}\times\mathcal{X}\to[0,1]$ that takes the T1 and T2 MRI volumes to optimize a standard binary cross-entropy objective function. This trained classifier, together with $\mathcal{D}_{T}$ will then be used for the filtering process of multi-rater cleaning technique CrowdLab [12], denoted by $z:[0,1]\times\mathcal{Y}\to\{0,1\}$, which generates a pseudo clean label for each sample. Formally, CrowdLab’s pseudo labeling process is defined by:
@@ 3.3 Multi-modal Human-AI Collaborative Classification
Given the pseudo clean labels $y^{(i)}$ from (3), the dataset for the final model training is defined as:
@@
The pre-trained encoder from (1), parameterized by $\phi^{*}$, is utilized to initialize the T1 and T2 MRI feature extractors, respectively defined by $g_{\hat{\phi}}:\mathcal{X}\to\mathcal{F}$ and $g_{\phi^{\prime}}:\mathcal{X}\to\mathcal{F}$, for the classification task. We also need to define a new feature extractor for processing the manual labels with a learnable module defined as $g_{\gamma}:\mathcal{Y}\to\mathcal{F}$. Such manual labels are used in the human-AI collaborative module. Hence, these extractors produce the T1, T2 and rater features as:
$\begin{split}\mathbf{F}_{t1}^{(i)}&=g_{\hat{\phi}}(\mathbf{x}_{t1}^{(i)}),\\ \mathbf{F}_{t2}^{(i)}&=g_{\phi^{\prime}}(\mathbf{x}_{t2}^{(i)}),\\ \mathbf{F}_{r}^{(i)}&=g_{\gamma}(\mathbf{y}^{(i)}).\end{split}$ (5)
@@
where $\mathbf{p}^{(i)}\in[0,1]^{2}$ denotes the probabilistic prediction, $\oplus$ represents the concatenation operator, and $\sigma$ is the softmax function.
Finally, the training of the whole model is performed by minimizing the binary cross-entropy loss, as follows:
@@
where the $p^{(i)}$ is the predicted probability of the positive class of $i$-th sample (i.e., the $2^{nd}$ dimension $\mathbf{p}^{(i)}$ in (6)), and $y^{(i)}$ is the pseudo clean label in $\mathcal{D}^{\prime}_{T}$. The goal is to learn the optimal parameters $\hat{\phi}^{*},\phi^{\prime*},\gamma^{*}$, and $\eta^{*}$ that minimize this loss across all $N$ samples.
@@ 4.1 Endometriosis Dataset
We first introduce our multi-modal multi-rater dataset annotated with imaging and surgery-based POD obliteration labels for the diagnosis of endometriosis. For the pre-training stage, we collected 5,867 unlabeled T1 MRI volumes and 8,984 unlabeled T2 MRI volumes from patients aged between 18 and 45 years old, where the volumes show female pelvis scans obtained from various MRI machines with varying resolutions. The pre-training pelvic MRI scans were obtained from the South Australian Medical Imaging (SAMI) service. For this paper, we included only pelvic MRI scans of female patients acquired using either T1- or T2-weighted imaging sequences. These scans were collected from multiple centers using various equipments, including the 1.5T Siemens Aera, 3T Siemens TrioTim, 1.5T Philips Achieva, 1.5T Philips Ingenia, and 1.5T Philips Intera. The T1 and T2 datasets include a wide variety of sequences, where each scan is acquired using a combination of the following parameters: spin echo sequences (e.g., Turbo Spin Echo, TSE) or gradient echo sequences (e.g., Volumetric Interpolated Breath-hold Examination, VIBE); orientation in either 2D planes (axial, sagittal, coronal) or 3D sequences (e.g., SPC); with or without fat saturation/suppression techniques (e.g., DIXON, SPIR, SPAIR); and with or without gadolinium contrast enhancement. We show the statistics of the size distribution information of the pre-training dataset in Table 1. As described in Section 3.1, this dataset is used as containing 14,851 independent scans for the Masked Autoencoder (MAE) self-supervised learning. To standardize the data, the volumes were resampled to $1\text{mm}\times 1\text{mm}\times 3\text{mm}$ voxels. Also, we apply 3D contrast-limited adaptive histogram equalization (CLAHE) to enhance image local contrast and refine edge definitions. We choose CLAHE, because it produced the best classification results among the different data pre-processing methods tested on an internal validation process (e.g., voxel value truncate.).
Table 1: Statistics of the volume size distribution of the pre-training dataset.
(a) T1 pre-training dataset
Dim Mean Std Max Min
0 326.67 93.73 550 66
1 271.97 88.78 530 60
2 79.64 24.5 430 50
(b) T2 pre-training dataset
Dim Mean Std Max Min
0 245.78 113.33 530 60
1 208.47 93.87 504 50
2 77.01 23.99 430 50
For the training of the human-AI collaborative POD obliteration classifier, we collected 82 pairs of T1/T2 MRI volumes with patients aged 18 to 45 years old. These scans were obtained across multiple clinical sites, with each case annotated by three experienced clinicians who work in clinics specialized in the imaging-based diagnosis of endometriosis. The scans were acquired using a standardized protocol for endometriosis examination and they show a specific region surrounding the uterus, which is the area where the signs of POD obliteration are more visible. MRIs from patients with a history of hysterectomy or those with large pelvic lesions that limited adequate assessment of the POD were excluded. The scanners are of the following models: SIEMENS Aera, SIEMENS Espree, SIEMENS MAGNETOM Sola, and SIEMENS MAGNETOM Vida. The volume types are: T1-weighted MRI images, obtained using a Volume Interpolated Breath-hold Examination technique with Dixon fat-water separation, acquired in the transverse plane; and T2-weighted MRI images, acquired using the SPACE sequence, characterized by non-saturation techniques, taken in the coronal plane, and with isotropic voxel dimensions. All 3D volumes were reoriented to a Right-Anterior-Superior (RAS) coordinate system to standardize anatomical alignment, then resampled with an output spacing. Similar to the pre-training, to standardize the data, the volumes were resampled to $1\text{mm}\times 1\text{mm}\times 3\text{mm}$ voxels. We show the statistics of the size distribution information of the POD obliteration dataset in Table 2
Table 2: Statistics of the volume size distribution of the POD obliteration dataset.
(a) T1 modality in the dataset
Dim Mean Std Max Min
Depth 383.91 91.88 576 260
Height 312.10 72.61 464 250
Width 72.16 8.46 80 51
(b) T2 modality in the dataset
Dim Mean Std Max Min
Depth 389.80 18.40 448 384
Height 192.30 26.22 336 160
Width 289.71 19.37 354 192
This training set with 82 volumes is subdivided into 62 volumes for training and 20 for validation that are used for cross validation for hyper-parameter selection. We further collect 30 cases that contain ground truth annotation of POD obliteration from surgical reports. These cases serve as gold standards for testing. We show in Table 6 the distribution of normal vs abnormal (i.e., POD obliteration) images in the training and testing sets. Note that while in the testing set, we have ground truth from surgery, in the training set, we only have the labelers’ annotation. We also use CLAHE pre-processing in this dataset. Similar to the pre-training, to standardize the data, the volumes were resampled to $1\text{mm}\times 1\text{mm}\times 3\text{mm}$ voxels.
Table 3: Number of normal and abnormal (i.e., with POD obliteration) images in the training and testing sets. In the training set, ‘A1’ to ‘A3’ are the three annotators, ‘MV’ represents the annotations from majority voting, and ‘CL’ denotes the annotations from CrowdLab [55]. In the testing set, we have the grount truth (GT) from surgery.
POD Training Set Testing Set
A1 A2 A3 MV CL GT
Normal 62 48 36 51 47 15
Abnormal 20 34 46 31 34 15
@@ 4.2 Implementation Details
For model pre-training, the input volumes are cropped and possibly zero-padded to achieve the dimension of 64 × 128 × 128 voxels. To maintain consistency with the pre-training dataset, the endometriosis training and testing samples are manually centered at the uterine region which is the most important region for the detection of POD obliteration. Later we center crop each volume to the same dimensions as pre-training data, i.e. $64\times 128\times 128$ voxels, for training and testing. In both pre-training and the training of the human-AI collaborative POD obliteration classifier, the multi-modal encoder for each modality is a transformer with 12 blocks. The majority vote classifier has a 3D-ResNet50 as its backbone network [58]. For the human-AI collaborative POD obliteration classifier training, we use 5 epochs for model optimization warming up. AdamW optimizer and base learning rate of 1e-3 with cosine annealing [59] learning rate tuning strategy are adopted. Three multi-rater labels from three different annotators are incorporated into the training process. In the testing phase, the scans are also cropped in the uterine regions and the clinical surgical results serve as the ground truth for accurate evaluation. Using a cross-validation procedure, with a training set containing 62 volumes and a validation set comprising 20 volumes, we observe signs of overfitting after 60 epochs. To prevent this, we implemented early stopping and halted the training at the 60th epoch. Note that the majority voting is only produced for the consensus pseudo label required for the training of the model $f_{\theta}(.)$ to be used by CrowdLab, as explained in Eq. 3. Once the pseudo clean labels are generated by CrowdLab, the majority voting will no longer be needed. All experiments in the paper are run on an NVIDIA GeForce RTX 3090. The main Python libraries used in our implementation are: torch 1.7.1+cu110, torchvision 0.8.2+cu110, scikit-image and scipy.
@@ 4.3 Quantitative Evaluation Settings
We also show an ablation study that compares the performance of each annotator without using the system, with the performance of different combinations of annotators to be used in the human-AI collaboration, and the performance of the proposed HAICOMM.
5 Results and Discussion
5.1 Model Performance
Table 4: Performance comparison of HAICOMM and other models on the test set in terms of accuracy and AUROC (and respective improvements of HAICOMM) with the testing ground-truth labels from surgical reports. The standard deviation is calculated by inference time bootstrapping. “Majority Vote” denotes a purely manual classification using the majority vote of the three annotators. SSR and ProMix w/ “GT1”, “GT2” and “GT3” mean noisy-label learning models trained with labels from annotators #1, #2 and #3, respectively. “CL” denotes noisy-label learning models trained with labels $y^{(i)}$ from Eq. (3) produced by the multi-rater learning method CrowdLab. “HAIC” represents a model that collaborates with the annotators and that is trained with labels from CrowdLab. The best results for each column are in bold.
@@ 5.1 Model Performance
Human Majority Vote 0.70±0.00 14.29% - -
SSR w/ GT1 0.58±0.02 37.93% 0.58±0.01 53.83%
Noisy SSR w/ GT2 0.60±0.04 33.33% 0.58±0.01 51.67%
Label SSR w/ GT3 0.57±0.02 41.17% 0.56±0.06 57.07%
Learning ProMix w/ GT1 0.53±0.02 50.01% 0.54±0.04 64.75%
ProMix w/ GT2 0.62±0.03 29.72% 0.58±0.02 52.63%
ProMix w/ GT3 0.52±0.05 54.83% 0.57±0.02 56.54%
Multi- SSR w/ CL 0.62±0.02 29.72% 0.59±0.01 50.82%
rater ProMix w/ CL 0.65±0.04 23.08% 0.54±0.03 63.32%
HAIC SSR w/ HAIC 0.68±0.08 17.08% 0.74±0.01 19.37%
ProMix w/ HAIC 0.67±0.06 20.00% 0.74±0.04 19.98%
Ours HAICOMM 0.80±0.04 - 0.89±0.06 -
The performance results in Table 4 show that the proposed HAICOMM outperforms other competing models by a large margin across the accuracy and AUC measures. Relative improvements vary from 9.10% to 54.83% on accuracy and 19.37% and 64.75% on AUROC. The standard deviation is calculated by inference time bootstrapping. Also, we provide the the ROC Curve of HAICOMM models and its counterparts SSR and ProMix with CrowdLab labels in Fig. 3.
Figure 3: The ROC Curve of HAICOMM models and its counterparts.